OpenStudy (anonymous):

what is this in simplest form... (x/y^2)/(x^2/y^5)

OpenStudy (anonymous):

\[\frac{x\times y^5}{y^2\times x^2}\] is a start, then cancel

OpenStudy (anonymous):

so is it y^3/x

OpenStudy (anonymous):

yup